OpenStudy (anonymous):

Tazewell high school decided to take their students on a field trip to the movie theater. A total of 250 people went on the trip. Adults pay $4.50 for a ticket and students pay $2.50 for a ticket. How many students and how many adults went to the movies if they paid a total of $805 at the movie theater?

OpenStudy (anonymous):

..?

OpenStudy (radar):

Let x = the number of students then 250-x = the number of adults. As they are a total of 250 people attending. Now for the money part, we need to figure how much each group paid because we know the total amount paid was $805. We also know the price of each ticket. 2.50x + 4.50(250-x) = $805 look this equation over very carefully, be sure you know where each term came from. 2.50x is the amount the students paid and the 4.50(250-x) is the amount the adults paid. Lets add them and get the $805 total. 2.50x + 1125 -4.50x = $805 -2x =$805-$1125= -320 -2x=-$320 x=160 students 250-x = 250-160=90 adults.

OpenStudy (radar):

Did you follow with understanding?

OpenStudy (anonymous):

yes thank you soo much Im writing notes on it...so i can try to do it on my own next time.

OpenStudy (radar):

Very good, there will be more problems similar to that, some about coins and their sum etc. But they can be worked out like the above. Good luck with them.

OpenStudy (anonymous):

okay thank you!